# Understanding the Effectiveness of Large Language Models in Detecting Security Vulnerabilities

### v2

@@ 3.5. RQ5: LLMs vs Static Analysis Tools
Experiment setup. We next explore how GPT-4 (CWE-DF) compares against CodeQL. Since CodeQL requires building projects before analysis and the real-world datasets contain large projects, we limit our focus to the three synthetic datasets, namely OWASP and Juliet Java and Juliet C/C++. In this study, we run the official CodeQL queries designed for the top 25 CWEs. Table 6 presents results from CodeQL and GPT-4 on the three datasets. CodeQL performs better than GPT-4 (CWE-DF) in terms of accuracy on all three datasets. GPT-4 reports a 0.09 higher F1 than CodeQL on Juliet C/C++ while CodeQL reports a 0.03 and 0.18 higher F1 on OWASP and Juliet Java respectively.
| Dataset/Model | CodeQL | GPT-4 (CWE-DF) | ||||||
@@
| OWASP | 0.63 | 0.58 | 0.96 | 0.72 | 0.55 | 0.53 | 1.00 | 0.69 |
| Juliet Java | 0.94 | 0.92 | 0.95 | 0.94 | 0.70 | 0.63 | 0.98 | 0.76 |
| Juliet C/C++ | 0.72 | 0.99 | 0.44 | 0.61 | 0.59 | 0.55 | 0.98 | 0.70 |
@@
3.6. RQ6: LLMs vs Deep-Learning-Based Tools
We compare LLMs against two prior deep learning-based approaches: 1) DeepDFA (Steenhoek et al., 2024), which trains Graph Neural Networks using embeddings of control flow graphs and associated data flow facts, and 2) LineVul (Fu and Tantithamthavorn, 2022), which is a transformed-based model trained using token-based representation of code.
Experiment setup. We used the DeepDFA and LineVul versions from DeepDFA’s latest artifact version (DeepDFA artifact, 2024). For CVEFixes C/C++ and Juliet C/C++, we use the same test set as our main evaluation §3.1. We split the remaining dataset into training (80%) and validation (20%) sets. We also compare the results with the BigVul (Fan et al., 2020) dataset, used in the prior work. Because BigVul is a much larger real-world dataset, it can potentially improve the performance of the tools. We explore two evaluation settings: 1) train and test on the same dataset and 2) train on one and test on a different dataset to evaluate the generalizability of these models. Table 7 presents the results. Each cell reports the average results across three runs.
DL-based approaches have limited effectiveness on real-world datasets. We observe that both DeepDFA and LineVul obtain a maximum F1 score of 0.62 on CVEFixes C/C++, which is quite similar to the scores of GPT-4. For DeepDFA, while training on larger BigVul and Juliet C/C++ datasets improves its F1 scores, the accuracy scores remain close to 50%. On the other hand, LineVul is able to learn better even when using CVEFixes C/C++, which is relatively smaller. Interestingly, when trained on BigVul, LineVul obtains a poor F1 score of 0.02 on CVEFixes C/C++, which might indicate an over-fitting problem. We observe a similar pattern when LineVul is tested on Juliet C/C++ using CVEFixes C/C++ or BigVul training datasets. For Juliet C/C++, both tools obtain better scores compared to GPT-4. LineVul obtains a perfect score indicating that Juliet C/C++ has limited code patterns that are likely easier to learn. However, as we explain later, LineVul does not generalize well.
Generalizability across datasets. We observe that, for LineVul, the F1 scores drop drastically when trained on BigVul and tested on CVEFixes C/C++ or Juliet C/C++. Compared to LineVul, DeepDFA generalizes better, obtaining higher accuracy and F1 scores on Juliet C/C++ when trained using other datasets.
Trade-offs. DeepDFA involves significant inference overhead, due to the CFG extraction and dataflow analysis steps. LLMs, however, can use the textual representation of code and can operate on incomplete/partial programs. The use of data-flow and control-flow information in DeepDFA is evidently useful. We made similar observations with LLMs when using the CWE-DF prompt. On the other hand, LineVul, like LLMs, can leverage natural language information but has a generalization problem. Finally, both DeepDFA and LineVul provide binary labels and line numbers that are difficult to interpret. LLMs can additionally provide explanations, which are useful for further debugging (as shown in prior sections).
@@ 3.6. RQ6: LLMs vs Deep-Learning-Based Tools
| GPT-4 | CWE-DF | Juliet C/C++ | 0.59 | 0.55 | 0.98 | 0.70 |
| DeepDFA | Juliet C/C++ | Juliet C/C++ | 0.77 | 0.74 | 0.82 | 0.78 |
| DeepDFA | CVEFixes C/C++ | Juliet C/C++ | 0.64 | 0.72 | 0.46 | 0.55 |
| DeepDFA | BigVul | Juliet C/C++ | 0.74 | 0.70 | 0.82 | 0.77 |
| LineVul | Juliet C/C++ | Juliet C/C++ | 1.0 | 1.0 | 0.99 | 1.0 |
| LineVul | CVEFixes C/C++ | Juliet C/C++ | 0.51 | 0.62 | 0.14 | 0.22 |
| LineVul | BigVul | Juliet C/C++ | 0.43 | 0.42 | 0.39 | 0.41 |
4. Related Work
@@
Static analysis tools for vulnerability detection. Static analysis tools search for pre-defined vulnerability patterns in code. Tools such as FlawFinder (FlawFinder, 2023) and CppCheck (CPPCheck, 2023) use syntactic and simple semantic analysis techniques to find vulnerabilities in C++ code. More advanced tools like CodeQL (Avgustinov et al., 2016), Infer (Fb Infer, 2023), and CodeChecker (Code Checker, 2023) employ semantic analysis techniques and can detect vulnerabilities in multiple languages. Static analysis tools rely on manually crafted rules and precise specifications of code behavior, which is difficult to obtain automatically. In contrast, while LLMs cannot always reliably perform end-to-end reasoning over code, we observe that LLMs can automatically identify such specifications by leveraging statistically learned rules from training data. Further, due to their semantic understanding of natural language, LLMs can also perform more contextual reasoning. These abilities can potentially be leveraged to improve static analysis tools.
@@
Deep Learning-based vulnerability detection. Several works have focused on using Deep Learning techniques for vulnerability detection. Earlier works such as Devign (Zhou et al., 2019), Reveal (Chakraborty et al., 2020), LineVD (Hin et al., 2022) and IVDetect (Li et al., 2021) leveraged Graph Neural Networks (GNNs) for modeling dataflow graphs, control flow graphs, abstract syntax trees and program dependency graphs. Other works explored alternate model architectures: VulDeePecker (Li et al., 2020) and SySeVR (Li et al., 2018) used LSTM-based models on slices and data dependencies while Draper used Convolutional Neural Networks. Recent works demonstrate that Transformer-based models fine-tuned on the task of vulnerability detection can outperform specialized techniques (CodeBERT, LineVul (Fu and Tantithamthavorn, 2022), UnixCoder). DeepDFA (Steenhoek et al., 2024) and ContraFlow (Cheng et al., 2022) learn specialized embeddings that can further improve the performance of Transformer-based vulnerability detection tools. To the best of our knowledge, these techniques provide binary labels for vulnerability detection and cannot classify the type of vulnerability. Thapa et al. (2022) explore whether Language Models fine-tuned on multi-class classification can perform well where the classes correspond to groups of similar types of vulnerabilities. In contrast, we study some of the largest Language Models, such as GPT-4, and perform a much granular CWE-level classification, generate human-readable informal specifications and explore various prompting techniques that allow using the LLMs out-of-the-box.
LLMs for automated software engineering. Several recent approaches have demonstrated that LLMs can be effectively leveraged to improve the state-of-the-art performance in various traditional software engineering tasks such as automated program repair (Xia et al., 2023; Joshi et al., 2023; Xia and Zhang, 2022), test generation (Lemieux et al., 2023; Deng et al., 2023), code evolution (Zhang et al., 2023), and fault localization (Yang et al., 2024). However, unlike these approaches, we find that LLMs have limited vulnerability detection abilities.
Recently, Li et al. (Li et al., 2024) developed Llift, an approach that combines LLMs with static analysis to detect Use Before Initialization (UBI) bugs in Linux kernel. While they focus on a specific class of bugs, their approach supports the observations we make in Section 3.5 on the complementary nature of static analysis and LLMs. To the best of our knowledge, our work is the first comprehensive and general study of vulnerability detection abilities of LLMs across a broad range of vulnerabilities (25 CWEs) across two languages.
5. Threats to Validity
@@
External. The choice of LLMs and datasets may bias our evaluation and insights. To address this threat, we choose multiple popular synthetic and real-world datasets across two languages: Java and C++. We also choose both state-of-the-art closed-source and open-source LLMs. However, our insights may not generalize to other languages or datasets not studied in this paper.
@@
Internal. Owing to the non-deterministic nature of LLMs and single experiment runs per benchmark, our observations may be biased. To mitigate this threat, we use a temperature of zero to ensure determinism across all LLMs. While this works well for locally run CodeLlama models, it is well-known that GPT-4 and GPT-3.5 might still return non-deterministic results. However, due to the large number of benchmarks we evaluate, the non-determinism should balance out across the datasets. Further, given the poor effectiveness of all LLMs across the board, we do not expect our results to significantly change with re-runs.
@@
Our evaluation code and scripts may have bugs, which might bias our results. Our manual analysis of results may lead to erroneous inferences. To address this threat, multiple co-authors reviewed the code regularly and actively fixed issues. Further, multiple co-authors of the paper independently analyzed the results and discussed them together to mitigate any discrepancies.
6. Conclusion
@@
In this work, we performed a comprehensive analysis of LLMs for security vulnerability detection. Our study reveals that both closed-source LLMs, such as GPT-4, and open-source LLMs, like CodeLlama, perform modestly at vulnerability detection for both Java and C/C++. Their performance on vulnerability detection is lower when dealing with real-world code. However, we find that even in cases where the models produce incorrect predictions, they identify relevant sources, sinks and sanitizers for dataflow analysis. Hence, we believe that an interesting future direction is to develop neuro-symbolic techniques that combine the intuitive reasoning abilities of LLMs with symbolic tools such as logical reasoning engines and static code analyzers for more effective and interpretable solutions.
Appendix A Appendix
A.1. Dataset Processing and Selection
We perform a data processing and cleaning step for each dataset before evaluating them with LLMs.
OWASP. We remove or anonymize information in OWASP benchmarks that may provide obvious hints about the vulnerability in a file. For instance, we change package, variable names, and strings such as “owasp”, “testcode”, and “/sqli-06/BenchmarkTest02732” to other pre-selected un-identifying names such as “pcks”, “csdr”, etc. We remove all comments in the file because they may explicitly highlight the vulnerable line of code or may have irrelevant text (such as copyright info), which may leak information. These changes, however, do not change the semantics of the code snippets.
Juliet Java and C/C++. Similar to OWASP, we remove all comments and transform all identifiers that leak identifying information in all test cases in the Juliet benchmark. For instance, we change “class CWE80_XSS_CWE182_Servlet_connect_tcp_01” to “class MyClass”. The Juliet benchmark provides the vulnerable (named as “bad”) and non-vulnerable (named as “good*”) methods in the same file. For easier evaluation, we perform a pre-processing step to split each file into two, each containing either a vulnerable or non-vulnerable method. Juliet also contains special benchmarks that have dependencies across multiple (2-5) files. We skip these benchmarks because they are typically too big to fit into the LLM prompt. Hence, the number of test cases after the data processing step in Juliet is reduced (as shown in Table 4).
CVEFixes. For each CVE, CVEFixes provides the methods that were involved in the fix commit. It also includes the method code in the parent commit, i.e., the method version before the fix. We collect all methods in the fix commit and the parent commit and label them as vulnerable and non-vulnerable, respectively. Similar to other datasets, we also remove all comments in the method code. While CVEFixes contains methods across multiple programming languages, we only collect C/C++ and Java methods for our study.
A.2. Prompting Strategies
The Basic prompt is presented in Listing 1, CWE specific prompt in Listing 2 and Dataflow analysis-based prompt in Listing 3.
A.3. Other Prompting Strategies
In addition to the prompting strategies presented in our main evaluation, we considered other popular prompting strategies such as Few-shot prompting and Chain-of-thought prompting in a limited experimental setting. For the few-shot prompt (CWE-Few-shot), we included two examples of the task (one with a vulnerability and one without) in the CWE specific prompt before providing the target code snippet. For the chain-of-thought prompt (CWE-CoT), we explicitly ask the model to provide a reasoning chain before the final answer by adding a “Let’s think step-by-step” statement at the end of the CWE specific prompt.
Table 8 and Table 9 present the results from GPT-4 with various prompting strategies on a random subset of 100 samples of the Juliet Java and CVEFixes C/C++ datasets respectively. The CWE-DF prompt reports the highest accuracy of 69% and the highest F1 score of 0.75 on the Juliet Java dataset. The CWE-DF prompt reports a 0.05 higher F1 score than the CWE-CoT prompt and a 0.03 higher F1 score than the CWE-Few-shot prompt. This difference is much more prominent on the CVEFixes C/C++ dataset where the CWE-DF prompt reports a 0.34 higher F1 score than the CWE-CoT prompt and a 0.31 higher F1 score than the CWE-Few-shot prompt. Moreover, the CWE-Few-shot prompt reported a 0.2 lower F1 score than the CWE specific prompt on the CVEFixes C/C++ dataset while requiring more tokens. Our analysis of the few-shot prompts suggests that providing more examples may not be a useful strategy for vulnerability detection. Because the potential set of vulnerable code patterns is quite large, the provided examples hardly make a difference to LLMs’ reasoning abilities. Hence, it may be more useful to use prompts that instead elicit reasoning or explanations of some kind before detecting if the given snippet is vulnerable. The CWE-CoT prompt, however, does not help with reasoning always, as it either performed at par or worse than the Dataflow analysis-based prompt.
| Model | Prompt | Metrics | ||||
| A | P | R | F1 | |||
| GPT-4 | CWE | 0.65 | 0.58 | 0.96 | 0.72 | |
| GPT-4 | CWE-Few-shot | 0.65 | 0.58 | 0.94 | 0.72 | |
| GPT-4 | CWE-CoT | 0.69 | 0.64 | 0.79 | 0.70 | |
| GPT-4 | CWE-DF | 0.69 | 0.61 | 0.96 | 0.75 | |
| Model | Prompt | Metrics | |||||
| A | P | R | F1 | ||||
| GPT-4 | CWE | 0.55 | 0.54 | 0.58 | 0.56 | ||
| GPT-4 | CWE-Few-shot | 0.49 | 0.38 | 0.34 | 0.36 | ||
| GPT-4 | CWE-CoT | 0.52 | 0.37 | 0.30 | 0.33 | ||
| GPT-4 | CWE-DF | 0.56 | 0.56 | 0.83 | 0.67 | ||
Learning from these experiments, we selected the CWE specific prompt, Dataflow analysis-based prompt, in addition to the Basic prompt, for our main evaluation with LLMs.
A.4. Detailed metrics across all LLMs and Datasets
Table 10 presents the metrics for all LLMs and Datasets across all prompts.
| Model | Prompt | OWASP | Juliet Java | CVEFixes Java | Juliet C/C++ | CVEFixes C/C++ | ||||||||||||||||||||
| A | P | R | F1 | A | P | R | F1 | A | P | R | F1 | A | P | R | F1 | A | P | R | F1 | |||||||
| GPT-4 | Basic | 0.52 | 0.51 | 1.00 | 0.68 | 0.56 | 0.54 | 0.85 | 0.66 | 0.50 | 0.50 | 0.34 | 0.41 | 0.54 | 0.52 | 0.92 | 0.67 | 0.51 | 0.50 | 0.57 | 0.54 | |||||
| GPT-4 | CWE | 0.54 | 0.52 | 1.00 | 0.69 | 0.69 | 0.63 | 0.97 | 0.76 | 0.55 | 0.56 | 0.44 | 0.50 | 0.58 | 0.54 | 0.95 | 0.69 | 0.52 | 0.52 | 0.52 | 0.52 | |||||
| GPT-4 | CWE-DF | 0.55 | 0.53 | 1.00 | 0.69 | 0.70 | 0.63 | 0.98 | 0.76 | 0.53 | 0.53 | 0.59 | 0.56 | 0.59 | 0.55 | 0.98 | 0.70 | 0.52 | 0.51 | 0.76 | 0.61 | |||||
| GPT-3.5 | Basic | 0.53 | 0.52 | 0.72 | 0.60 | 0.58 | 0.57 | 0.71 | 0.63 | 0.46 | 0.35 | 0.09 | 0.15 | 0.49 | 0.49 | 0.64 | 0.56 | 0.52 | 0.55 | 0.19 | 0.29 | |||||
| GPT-3.5 | CWE | 0.55 | 0.54 | 0.62 | 0.58 | 0.52 | 0.52 | 0.55 | 0.54 | 0.47 | 0.41 | 0.12 | 0.19 | 0.49 | 0.49 | 0.70 | 0.58 | 0.51 | 0.54 | 0.19 | 0.28 | |||||
| GPT-3.5 | CWE-DF | 0.51 | 0.51 | 0.93 | 0.66 | 0.40 | 0.44 | 0.73 | 0.55 | 0.54 | 0.53 | 0.66 | 0.59 | 0.40 | 0.44 | 0.77 | 0.56 | 0.52 | 0.52 | 0.75 | 0.61 | |||||
| CL-34B | Basic | 0.51 | 0.51 | 1.00 | 0.67 | 0.47 | 0.48 | 0.85 | 0.62 | 0.50 | 0.50 | 0.28 | 0.36 | 0.50 | 0.50 | 0.93 | 0.65 | 0.51 | 0.51 | 0.19 | 0.28 | |||||
| CL-34B | CWE | 0.57 | 0.54 | 0.94 | 0.69 | 0.49 | 0.49 | 0.94 | 0.65 | 0.50 | 0.51 | 0.16 | 0.25 | 0.53 | 0.52 | 0.98 | 0.68 | 0.51 | 0.54 | 0.08 | 0.14 | |||||
| CL-34B | CWE-DF | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | |||||
| CL-13B | Basic | 0.52 | 0.51 | 0.98 | 0.67 | 0.47 | 0.48 | 0.72 | 0.58 | 0.50 | 0.50 | 0.13 | 0.21 | 0.46 | 0.48 | 0.79 | 0.59 | 0.51 | 0.52 | 0.22 | 0.31 | |||||
| CL-13B | CWE | 0.52 | 0.51 | 0.98 | 0.67 | 0.50 | 0.50 | 0.89 | 0.64 | 0.48 | 0.47 | 0.29 | 0.36 | 0.53 | 0.51 | 0.98 | 0.67 | 0.52 | 0.52 | 0.56 | 0.54 | |||||
| CL-13B | CWE-DF | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 0.96 | 0.66 | |||||
| CL-7B | Basic | 0.56 | 0.54 | 0.81 | 0.65 | 0.77 | 0.79 | 0.72 | 0.75 | 0.47 | 0.46 | 0.42 | 0.44 | 0.62 | 0.58 | 0.86 | 0.69 | 0.47 | 0.35 | 0.07 | 0.11 | |||||
| CL-7B | CWE | 0.50 | 0.50 | 1.00 | 0.67 | 0.52 | 0.51 | 0.99 | 0.67 | 0.51 | 0.51 | 0.84 | 0.63 | 0.51 | 0.50 | 0.99 | 0.67 | 0.50 | 0.50 | 0.84 | 0.63 | |||||
| CL-7B | CWE-DF | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.50 | 0.50 | 1.00 | 0.67 | 0.51 | 0.50 | 0.98 | 0.66 | |||||
A.5. Adversarial Attack Examples
| Attack | Example |
| Dead-code Injection | ⬇ int obj = new Object(...); if (obj == null) { /* deadcode */ } |
| Variable Renaming | ⬇ String var3 = (String) names.nextElement(); String[] values = request.getParameterValues(var3); |
| Branch Insertion | ⬇ boolean var11 = true; if (var11) { configManager.init(); } |
Table 11 presents the three adversarial attacks that we implemented: dead-code injection, variable renaming, and branch insertion. Each attack contains multiple code patterns, which are randomly chosen when perturbing a given code snippet. We implement these attacks for Java and C/C++. For a given code snippet, we only apply an attack only once.
A.6. More Details on CodeQL vs. GPT-4
Table 12 presents the CWE-wise distribution of CodeQL results. CodeQL shows relatively better performance on Juliet Java (with an F1 score over 0.75 across CWEs) than Juliet C/C++ (with two CWEs reporting F1 scores under 0.4). On Juliet Java, CodeQL re- ports the highest performance on CWE-78 with an F1 score of 0.92. Interestingly, CWE-78 reports the worst results on the other two datasets (F1 score of 0.67 with OWASP and 0.03 with Juliet C/C++).
[fontsize=,breaklines,escapeinside=——,bgcolor=anti-flashwhite]cexamples/juliet_cpp_47137_highlighted.txt
In Listing LABEL:fig:julietcpp47137, the model incorrectly reasons about the sanitization by overlooking the & which would cause both the expressions to be evaluated. The identified sources and sinks are correct, however.
| Dataset | CWE | CodeQL | GPT-4 (CWE-DF) | ||||||
| A | P | R | F1 | A | P | R | F1 | ||
| OWASP | 22 | 0.52 | 0.50 | 1.00 | 0.67 | 0.48 | 0.48 | 1.00 | 0.64 |
| OWASP | 78 | 0.56 | 0.53 | 0.77 | 0.63 | 0.48 | 0.48 | 1.00 | 0.65 |
| OWASP | 89 | 0.57 | 0.54 | 1.00 | 0.70 | 0.52 | 0.52 | 1.00 | 0.68 |
| OWASP | 79 | 0.79 | 0.71 | 1.00 | 0.83 | 0.65 | 0.59 | 1.00 | 0.74 |
| Juliet Java | 78 | 0.94 | 0.86 | 1.00 | 0.92 | 0.57 | 0.47 | 1.00 | 0.64 |
| Juliet Java | 89 | 0.88 | 0.85 | 1.00 | 0.92 | 0.69 | 0.68 | 1.00 | 0.81 |
| Juliet Java | 476 | 0.76 | 0.76 | 0.76 | 0.76 | 0.79 | 0.75 | 0.88 | 0.81 |
| Juliet Java | 190 | 0.97 | 1.00 | 0.92 | 0.96 | 0.71 | 0.62 | 0.97 | 0.75 |
| Juliet C/C++ | 78 | 0.52 | 1.00 | 0.04 | 0.08 | 0.50 | 0.50 | 1.00 | 0.67 |
| Juliet C/C++ | 476 | 0.80 | 0.95 | 0.75 | 0.84 | 0.71 | 0.75 | 0.86 | 0.80 |
| Juliet C/C++ | 416 | 0.63 | 1.00 | 0.26 | 0.41 | 0.52 | 0.51 | 0.97 | 0.67 |
| Juliet C/C++ | 190 | 0.93 | 0.99 | 0.87 | 0.93 | 0.68 | 0.60 | 0.98 | 0.74 |
A.7. Qualitative analysis of GPT-4 responses
We first present examples where the dataflow analysis from the CWE-DF prompt is useful. Consider the code snippet in Listing 4. In this snippet, the variable dir is indirectly being used to create a directory via the dirToCreate variable. GPT-4 correctly identifies that this path is not sanitized and could be used to create a directory in otherwise restricted locations. This could lead to CWE-22 (path traversal) as is rightly predicted by GPT-4. In Listing 5, the method write takes an array b and integers offset and length as inputs. The expression offset + length > b.length can lead to an Integer Overflow (CWE-190) if offset + length exceeds the maximum integer value and this is correctly identified by GPT-4. Finally, in Listing 6, there are multiple array accesses that could lead to CWE-125 (Out-of-bounds read). GPT-4 correctly identifies that the expression ciphertextOffset + length + index could exceed the size of the ciphertext array, leading to this vulnerability.
A.8. CWE-125 (Out-of-bounds Read in Java vs. C++)
Listing 6 presents an example of Out-of-bounds read vulnerability in CVEFixes Java which GPT-4 predicts correctly. Codellama-34B also correctly predicts that this snippet is vulnerable. Listing 7 presents an example from CVEFixes C/C++ where Out-of-bounds read vulnerability is not possible but this is not evident from context. Codellama-34B’s explanation seems plausible given that little is known about the pointers in question.

### v3

@@ III-D RQ4: LLMs vs Static Analysis Tools
Experiment setup. We next explore how the LLMs compare against CodeQL. Since CodeQL requires building projects before analysis and the real-world datasets contain large projects, we limit our focus to the three synthetic datasets, namely OWASP, Juliet Java and Juliet C/C++. We use the official CodeQL queries designed for the top 25 CWEs. Table VIII presents results from CodeQL and the best performing LLM on the three datasets: CodeLlama-13B on OWASP, Llama-3.1-70B on Juliet Java and Qwen-2.5-14B on Juliet C/C++.
Finding 4.1: CodeQL performs better than the LLMs in terms of accuracy on all three datasets. CodeQL reports 0.07 and 0.15 higher F1 than CodeLlama-13B on OWASP and Llama-3.1-70B on Juliet Java respectively. However, Qwen-2.5-14B reports a 0.11 higher F1 on Juliet C/C++.
Finding 4.2: LLMs perform better than CodeQL on certain CWEs. A CWE-wise comparison of LLMs with CodeQL in Figure 3 reveals that LLMs report higher accuracies on CWE-22 (Path Traversal), CWE-78 (OS Command Injection), CWE-476 (NULL Pointer Dereference), 416 (Use After Free) on at least 2 / 3 datasets while CodeQL performs better on CWE-190 (Integer Overflow) on 2 datasets. Concretely,
CodeQL performs better than the LLMs on CWE-190 (Integer Overflow) with 11% and 21% higher accuracies than Llama-3.1-70B on Juliet Java and Juliet C/C++ respectively.
On the other hand, DSCoder-7B performs better than CodeQL on CWE-476 (NULL Pointer Dereference) with 12% higher accuracy on Juliet Java and only 1% lower accuracy on Juliet C/C++. Moreover, 6 / 16 LLMs report accuracies higher than CodeQL on CWE-476 from Juliet Java.
While CodeQL reports an extremely high accuracy of 95% on CWE-78 (OS Command Injection) with Juliet Java, it is outperformed by CodeLlama-13B by 1% and Codestral-22B by 7% on OWASP and Juliet C/C++ respectively. Interestingly, 7 / 16 LLMs report higher accuracies (by upto 10%) at detecting CWE-22 (Path Traversal) on OWASP.
Similarly, 3 LLMs perform better on CWE-416 (Use After Free) from Juliet C/C++ with DSCoder-7B reporting a whopping 24% higher accuracy than CodeQL.
Finding 4.3: CodeQL’s worse performance on some CWEs can be attributed to the very specific manually-written queries which may not cover all possible scenarios of the vulnerability. For example, CodeQL only detects CWE-78 (OS Command Injection) in C/C++ snippets when there exist system commands that take a string of arguments (execl). This query cannot handle commands that take a list of arguments (eg., _execvp). Listing LABEL:fig:motivating-example provides an example of this scenario where CodeQL does not detect that the snippet is prone to OS Command Injection but GPT-4 (CWE-DF) correctly identifies _execvp as a vulnerable sink. Listing LABEL:fig:julietcpp43583 presents an example where CodeQL correctly predicts that the target snippet is vulnerable to Integer Overflow while GPT-4 with CWE-DF does not. However, the model correctly identifies the sources, sinks and even unsanitized dataflows in this case but fails to faithfully reason over them when predicting the vulnerability. These examples support the observation from Section III-B that LLMs can be used to identify sources and sinks relevant to the target vulnerability (which can be missed by manually written queries) while static analysis tools like CodeQL can be used to reason over them. More examples are presented in the Appendix E.
| Dataset/Model | CodeQL | Best performing LLM | ||||||
@@
| OWASP | 0.63 | 0.58 | 0.96 | 0.72 | 0.60 | 0.58 | 0.74 | 0.65 |
| Juliet Java | 0.94 | 0.92 | 0.95 | 0.94 | 0.76 | 0.71 | 0.89 | 0.79 |
| Juliet C/C++ | 0.72 | 0.99 | 0.44 | 0.61 | 0.65 | 0.60 | 0.89 | 0.72 |
@@
III-E RQ5: LLMs vs Deep Learning-Based Tools
We compare LLMs against two prior deep learning-based approaches: 1) DeepDFA [10], which trains Graph Neural Networks using embeddings of control flow graphs and associated data flow facts, and 2) LineVul [11], which is a transformed-based model trained using token-based representation of code.
Experiment setup. Our main aim in this experiment is to check the generalizability of these techniques on real-world datasets beyond the datasets they are trained on. We use CVEFixes C/C++ with the 1000 samples from our main evaluation §III-A as the real-world test set. We train on three different datasets:Juliet C/C++, CVEFixes C/C++ (excluding samples in the test set) and BigVul [34](the original C/C++ dataset that these models were trained on). We use an 80/20 train-validation split while training on these datasets. We use the DeepDFA and LineVul versions from DeepDFA’s latest artifact [42]. Table IX presents the results, averaged across three runs.
Finding 5.1: DL-based approaches have limited effectiveness on real-world datasets. DeepDFA and LineVul trained on CVEFixes C/C++ training set report accuracies of 51% and 59% on CVEFixes C/C++ test set respectively while the best performing LLM on this dataset, Qwen-2.5-32B, reports an accuracy of 56%. This is in stark contrast to the performance of these techniques on BigVul where they report accuracies higher than 98%. Moreover, Qwen-2.5-32B reports an F1 score of 0.65 which is 0.42 and 0.04 higher than DeepDFA and LineVul respectively.
Finding 5.2: DL-based approaches do not generalize across datasets. When trained on Juliet C/C++ or BigVul, both DeepDFA and LineVul report accuracies and F1 scores lower than Qwen-2.5-32B by upto 6% and 0.63 respectively.
Finding 5.3: LLMs are preferable over DL-based approaches due to low inference overheads and natural language explanations. DeepDFA involves significant inference overhead, due to the CFG extraction and dataflow analysis steps. LLMs, however, can use the textual representation of code and can operate on incomplete/partial programs. The use of data-flow and control-flow information in DeepDFA is evidently useful. We made similar observations with LLMs when using the CWE-DF prompt. On the other hand, LineVul, like LLMs, can leverage natural language information but has a generalization problem. Finally, both DeepDFA and LineVul provide binary labels and line numbers that are difficult to interpret. LLMs can additionally provide explanations, which are useful for further debugging (as shown in prior sections).
@@ III-E RQ5: LLMs vs Deep Learning-Based Tools
IV Related Work
@@
Static analysis tools for vulnerability detection. Tools such as FlawFinder [43] and CppCheck [44] use syntactic and simple semantic analysis techniques to find vulnerabilities in C++ code. More advanced tools like CodeQL [5], Infer [45], and CodeChecker [46] employ semantic analysis techniques and can detect vulnerabilities in multiple languages. Static analysis tools rely on manually crafted rules and precise specifications of code behavior, which is difficult to obtain automatically. In contrast, while LLMs cannot always reliably perform end-to-end reasoning over code, we find that they are capable of automatically identifying these specifications which can be leveraged to improve static analysis-based detection tools.
@@
Deep Learning-based vulnerability detection. Several works have focused on using Deep Learning techniques for vulnerability detection. Earlier works such as Devign [47], Reveal [48], LineVD [49] and IVDetect [50] leveraged Graph Neural Networks (GNNs) for modeling dataflow graphs, control flow graphs, abstract syntax trees and program dependency graphs. Other works explored alternate model architectures: VulDeePecker [51] and SySeVR [52] used LSTM-based models on slices and data dependencies while Draper used Convolutional Neural Networks. Recent works demonstrate that Transformer-based models fine-tuned on the task of vulnerability detection can outperform specialized techniques (CodeBERT, LineVul [11], UnixCoder). DeepDFA [10] and ContraFlow [53] learn specialized embeddings that can further improve the performance of Transformer-based vulnerability detection tools. These techniques, however, provide binary labels for vulnerability detection and do not provide natural language explanations.
LLMs for automated software engineering. Recent approaches have demonstrated the effectiveness of LLMs on software engineering tasks such as automated program repair [13, 14, 15], test generation [16, 17], code evolution [18], and fault localization [19]. However, unlike these approaches, we find that scaling LLMs to larger sizes does not improve vulnerability detection abilities. [54] explore whether Language Models fine-tuned on multi-class classification can perform well where the classes correspond to groups of similar types of vulnerabilities. In contrast, we study whether LLMs can perform a much granular CWE-level classification through prompting. Recent work combining LLMs with static analysis to detect Use Before Initialization (UBI) bugs in the Linux Kernel [55] supports our claims in Section III-D (but focuses on a specific class of bugs). There are other concurrent studies evaluating LLMs on vulnerability detection [24, 25, 26, 27, 28]. Table II provides a comparison against these studies. Section III-A corroborates the common finding from these studies that LLMs do not perform well on real-world datasets. However, to the best of our knowledge, our study is the first work that qualitatively identifies prompts and vulnerability classes where LLMs perform well (and often better than other tools). Moreover, our study focuses on a larger/different class of state-of-the-art LLMs, datasets, languages, and vulnerability classes.
V Threats to Validity
@@
External. The choice of LLMs and datasets may bias our evaluation and insights. To address this threat, we choose multiple popular synthetic and real-world datasets across two languages: Java and C++. We also choose both state-of-the-art closed-source and open-source LLMs. However, our insights may not generalize to other languages or datasets.
@@
Internal. Owing to the non-deterministic nature of LLMs and single experiment runs per benchmark, our observations may be biased. To mitigate this threat, we use a temperature of 0 to ensure determinism across all LLMs. While this works well for locally run CodeLlama models, it is well-known that GPT-4 and GPT-3.5 might still return non-deterministic results. However, this should balance out across datasets and over the large number of benchmarks we evaluate on. Further, given similar results across LLMs on real-world-datasets, we do not expect significant changes with re-runs.
@@
Our evaluation code and scripts may have bugs, which might bias our results. Our manual analysis of results may lead to erroneous inferences. To address this threat, multiple co-authors reviewed the code regularly and actively fixed issues. Further, multiple co-authors independently analyzed the results and discussed them together to mitigate any discrepancies.
VI Conclusion
@@
In this work, we performed a comprehensive analysis of LLMs for security vulnerability detection. Our study reveals that both closed-source LLMs, such as GPT-4, and open-source LLMs, like CodeLlama, perform modestly at vulnerability detection for both Java and C/C++. However, we find specific vulnerability classes where LLMs excel (often performing better than static analysis tools, such as CodeQL). Moreover, we find that even in cases where the models produce incorrect predictions, they are capable of identifying relevant sources, sinks and sanitizers that can be used for downstream dataflow analysis. Hence, we believe that an interesting future direction is to develop neuro-symbolic techniques that combine the intuitive reasoning abilities of LLMs with symbolic tools such as logical reasoning engines and static code analyzers for more effective and interpretable solutions.